y2 - 18y + 4 = 70


Line 1
y^2 - 18y = 66
Line 2
y^2 - 18y + 81 = 66
Line 3
(y - 9)^2 = 66
Line 4
y - 9 = ± √66
Line 5
y = 9 ± √66

Answer by Alan3354(69443)

y2 - 18y + 4 = 70


Line 1
y^2 - 18y = 66
Line 2
y^2 - 18y + 81 = 66 ****** Add 81 to each side.
===========================
Line 3
(y - 9)^2 = 66
Line 4
y - 9 = ± √66
Line 5
y = 9 ± √66